# Cert-SSB: Toward Certified Sample-Specific Backdoor Defense

### v1

@@ V-E The Analysis of Computational Complexity
In this section, we analyze the computational complexity of Cert-SSB under an experimental setup running Ubuntu 22.04, equipped with an Intel Xeon Silver 4214 CPU, a Tesla V100-PCIE-32GB GPU, and CUDA 12.0. We particularly focus on the computational costs of the noise optimization and storage-update-based certification processes.
@@
The Complexity of Noise Optimization. Let $N$, $\lambda$, $T$ denote the number of samples in the training set, the poisoning rate, and the number of stochastic gradient ascent (SGA) iterations used for noise optimization, respectively. Since Cert-SSB first poisons a small subset of selected samples and then performs $T$ rounds of SGA optimization on each poisoned sample to obtain the optimized noise, its computational complexity is $\mathcal{O}(N\cdot\lambda\cdot T)$. Furthermore, Cert-SSB supports parallel processing, as the optimization process for each sample is independent. In practice, Cert-SSB requires only one round of SGA iteration to obtain sufficiently optimized noise. For instance, on the CIFAR-10 dataset, computing optimized noise for a single sample takes approximately 2 seconds. Therefore, the additional computational overhead of our method in the noise optimization phase is acceptable.
The Complexity of Storage-update-based Certification. In this stage, the defender adopts a storage-update-based method to dynamically update the certification process for $n$ samples, ensuring that the certification regions of different samples do not overlap. The computational complexity of this step is $\mathcal{O}(n)$, and it also supports parallel processing. For example, in batch mode, executing the storage-based update certification method on 2000 testing samples takes only about 2 seconds. Arguably, this computational overhead is negligible compared to the improvement in certification performance.
VI Conclusion
@@
In this paper, we revisited existing randomized smoothing-based certified backdoor defense methods and revealed that using fixed noise for all samples led to suboptimal certification performance. To address this issue, we proposed a sample-specific certified backdoor defense method ($i.e.$, Cert-SSB), which employed stochastic gradient ascent to iteratively optimize sample-specific noise in order to maximize the certification radius. The optimized noise was then injected into the poisoned training set to retrain multiple smoothed models, whose predictions are aggregated to obtain the final robust prediction. Since existing certification methods typically assumed a fixed noise level and thus did not apply to our setting, we further introduced a storage-update-based certification approach to improve certification accuracy and reliability. Extensive experiments on multiple benchmark datasets demonstrated that Cert-SSB significantly outperformed existing methods in terms of certification performance. We hope this work inspires future exploration of how sample-specific noise relates to model decision boundaries for better personalized certification.

### v2

@@ V-F The Analysis of Computational Complexity
In this section, we analyze the computational complexity of Cert-SSBD under an experimental setup running Ubuntu 22.04, equipped with an Intel Xeon Silver 4214 CPU, a Tesla V100-PCIE-32GB GPU, and CUDA 12.0. We particularly focus on the computational costs of the noise optimization and storage-update-based certification processes. A more detailed runtime analysis is provided in Appendix -F.
@@
The Complexity of Noise Optimization. Let $N$ and $T$ denote the number of training samples and the number of stochastic gradient ascent (SGA) iterations used for noise optimization, respectively. Since Cert-SSBD performs $T$ rounds of SGA-based noise optimization for each training sample to obtain sample-specific noise scales, the overall complexity of the noise optimization phase is $\mathcal{O}(N\cdot T)$. Furthermore, Cert-SSBD supports parallel processing, as the optimization process for each sample is independent. For instance, with T=100T{=}100, the per-sample optimization time is approximately 0.01 seconds for MNIST and 0.05 seconds for CIFAR-10. Therefore, the additional computational overhead of our method in the noise optimization phase is acceptable.
The Complexity of Storage-update-based Certification. In this stage, the defender adopts a storage-update-based method to dynamically update the certification process, resolving potential overlaps among certification regions across different samples. For time complexity, let $p$ be the probability that the certified region of a new sample $x_{n+1}$ overlaps with an existing certification region, and let $c$ denote the cost of computing a single certification region. The expected running time is $\mathcal{O}(np+(1-p)(2n+c))$. Specifically, $n$ comparisons are required to check whether the certified region of the new sample overlaps with an existing region; if no overlap occurs, a new region is computed (cost $c$) and the storage is updated (cost $n$). In practice, when overlaps are rare ($i.e.$, $p\approx 0$), the complexity simplifies to $\mathcal{O}(2n+c)$; our experiments confirm that this condition holds across all datasets. For instance, on MNIST and CIFAR-10, executing storage-update-based certification on the entire testing set takes only approximately 5 seconds, which is negligible compared to the certified performance gains. Besides, storing $n$ certified triplets incurs $\mathcal{O}(n)$ memory overhead in practice.
VI Potential Limitations and Future Directions
As one of the early studies on certified backdoor defenses, we admit that our method may exhibit certain potential limitations, which warrant further investigation in future work.
Firstly, Cert-SSBD introduces additional computational overhead compared to standard randomized-smoothing-based methods. This overhead primarily arises from two components: sample-specific noise optimization and storage-update-based certification. Specifically, the noise optimization procedure is a one-time offline preprocessing step that can be efficiently parallelized and therefore does not affect deployment-time inference efficiency. The storage-update-based certification mechanism only triggers updates when potential conflicts are detected, resulting in limited overhead in practice. As detailed in Section V-F, the overall computational cost remains controllable relative to the achieved certified performance gains. Nevertheless, future work may further explore strategies to accelerate and streamline our approach.
Secondly, Cert-SSBD maintains additional storage to support storage-update-based certification, which incurs extra memory and storage overhead. In practice, this overhead is typically manageable and primarily functions as a conservative safeguard to ensure certification consistency. In future work, we will investigate more compact storage representations as well as more efficient update and retrieval mechanisms to further reduce storage costs and improve scalability.
Thirdly, Cert-SSBD is currently evaluated primarily on image classification tasks and has not yet been systematically extended to other settings, such as text, speech, multimodal learning, or generative models. Arguably, the underlying principles of our approach are broadly applicable. In future work, we plan to extend sample-specific noise learning and the consistency-based certification mechanism to a wider range of modalities and task settings, and to evaluate their effectiveness under more complex attack scenarios and data distributions.
Finally, Cert-SSBD currently adopts isotropic scalar noise and adapts only the noise magnitude on a per-sample basis, without explicitly modeling direction-dependent decision-boundary geometry. This design choice is consistent with common practice in randomized-smoothing-based certification frameworks (see [45]) and does not affect our main conclusions regarding sample-specific noise learning and consistency-based certification. In future work, we plan to explore sample-specific anisotropic noise ($e.g.$, ellipsoidal certification schemes) to more accurately characterize local decision-boundary geometry.
VII Conclusion
@@
In this paper, we revisited existing randomized smoothing-based certified backdoor defense methods and revealed that using fixed noise for all samples led to suboptimal certification performance. To address this issue, we proposed a sample-specific certified backdoor defense method ($i.e.$, Cert-SSBD), which employed stochastic gradient ascent to iteratively optimize sample-specific noise in order to maximize the certification radius. The optimized noise was then injected into the poisoned training set to retrain multiple smoothed models, whose predictions are aggregated to obtain the final robust prediction. Since existing certification methods typically assumed a fixed noise level and thus did not apply to our setting, we further introduced a storage-update-based certification approach to improve certification accuracy and reliability. Extensive experiments on multiple benchmark datasets demonstrated that Cert-SSBD significantly outperformed existing methods in terms of certification performance. We hope this work inspires future exploration of how sample-specific noise relates to model decision boundaries for better personalized certification.